Analyzing the Growth in Shareholders’ Equity (Easy) The following numbers were calculated from the financial statements for a firm for 2012 and 2011:

  (2012) (2011)

Return on common equity (ROCE) 15.2% 13.3%

Return on net operating assets (RNOA) 11.28%    12.75%

Sales (millions) $ 16754 $ 11035

Average net operating assets (millions) $6981 $4414

Average net financial obligations (millions) $2225 $ 241

Average common equity (millions) $4756 $ 4,173

Explain to what extent the change in common equity from 2011 to 2012 is due to sales growth, net assets required to support sales, and borrowing.
Answer #1

Explain to what extent the change in common equity from 2011 to 2012 is due to sales growth, net assets required to support sales, and borrowing.

Answer:-

Change in CSE = 583

Change in sales = 5,719

Change in 1/ATO = 1/2.4 – 1/2.5 = 0.4167 – 0.4 = 0.0167

Change in NFO = 1,984

Change in CSE = 583 = (5,719 x 0.4) + (0.0167 x 16,754) – 1,984

= 2287.6 + 279.8 – 1,984.0

      ↓   ↓ ↓

    Due to Due to Due to

    Sales   NOA   Borrowing